# Editable Concept Bottleneck Models

### v1

@@ A.1 EK-FAC for CBMs
Besides, [14] proposed a new method that corrects the error in equation 14 which sets the $i$-th diagonal element of $\Lambda_{\Omega_{l-1}}\otimes\Lambda_{\Gamma_{l}}$ as
@@
Appendix B Proof of Concept-label-level Influence
@@
Because of equation 15, the first term of equation B equals 00. Then we have
@@
where $H_{\hat{f}}=\nabla^{2}_{\hat{f}}\sum_{i=1}^{n}L_{Y_{i}}(\hat{f},\hat{g})$ is the Hessian matrix of the loss function respect to $\hat{f}$, $L_{Y_{i}}(\hat{f},\hat{g})\triangleq L_{Y}(\hat{f}(\hat{g}(x_{i})),y_{i})$, and $\bar{g}_{e}$ is given in Theorem B.1.
@@
Appendix C Proof of Concept-level Influence
@@
The main ideas are as follows: (i) First, we define a new predictor $\hat{g}^{*}_{p_{M}}$, which has the same dimension as $\hat{g}$ and the same output as $\hat{g}_{-p_{M}}$. Then deduce an approximation for $\hat{g}^{*}_{p_{M}}$. (ii) Then, we consider setting $p_{r}=0$ instead of removing it, we get $\hat{f}_{p_{M}=0}$, which is equivalent to $\hat{f}_{-p_{M}}$ according to lemma C.1. We estimate this new predictor as a substitute. (iii) Next, we assume we only use the updated concept predictor $\hat{g}_{p_{M}}^{*}$ for one data $(x_{i_{r}},y_{i_{r}},c_{i_{r}})$ and obtain a new label predictor $\hat{f}_{ir}$, and obtain a one-step Newtonian iterative approximation of $\hat{f}_{ir}$ with respect to $\hat{f}$. (iv) Finally, we repeat the above process for all data points and combine the estimate of $\hat{g}$ in Theorem C.2, we obtain a closed-form solution of the influence function for $\hat{f}$.
@@
Now, we come to the approximation of $\hat{f}_{-p_{M}}$. Noticing that the input dimension of $f$ decreases to $k-|M|$. We consider setting $p_{r}=0$ for all data points in the training phase of the label predictor and get another optimal model $\hat{f}_{p_{M}=0}$. From lemma C.1, we know that for the same input $x$, $\hat{f}_{p_{M}=0}(x)=\hat{f}_{-p_{M}}$. And the values of the corresponding parameters in $\hat{f}_{p_{M}=0}$ and $\hat{f}_{-p_{M}}$ are equal.
@@
Theorem C.2 gives us the edited version of $\hat{g}^{*}_{-p_{M}}$. Substitute it into equation 25, and we get the final closed-form edited label predictor under concept level:
@@
Appendix D Proof of Data-level Influence
@@
The edited version of $\hat{g}_{-z_{G}}$ has been deduced as $\bar{g}_{-z_{G}}$ in theorem D.1, substituting this approximation into (37), then we have
@@
Combining (34) and (39), then we deduce the final closed-form edited label predictor as
@@ Appendix E Algorithm
3:  Use EK-FAC method to accelerate iHVP problem for $\hat{g}$
4:  for the $l$-th convolution layer of $\hat{g}$do
5:     Define the input activations $\{a_{j,t}\}$, weights $W=\left(w_{i,j,\delta}\right)$, and biases $b=\left(b_{i}\right)$ of this layer;
6:     Obtain the expanded activations $\llbracket{{A}}_{l-1}\rrbracket$ as:
7:     Compute the pre-activations:
8:     During the backpropagation process, obtain the $\mathcal{D}s_{i,t}$ as:
9:     Compute $\hat{\Omega}_{l-1}$ and $\hat{\Gamma_{l}}$:
10:     Perform eigenvalue decomposition of $\hat{\Omega}_{l-1}$ and $\hat{\Gamma}_{l}$, obtain $Q_{\Omega},{\Lambda}_{{\Omega}},{Q}_{\Gamma},{\Lambda}_{\Gamma}$, which satisfies
$\hat{\Omega}_{l-1}$ $=Q_{\Omega}{\Lambda}_{{\Omega}}{Q}_{{\Omega}}^{\top}$
$\hat{\Gamma}_{l}$ $={Q}_{\Gamma}{\Lambda}_{\Gamma}{Q}_{\Gamma}^{\top}$
11:     Define a diagonal matrix $\Lambda$ and compute the diagonal element as
12:     Compute $\hat{H}_{l}^{-1}$ as
13:  end for
14:  Splice $H_{l}$ sequentially into large diagonal matrices
15:  Update concept predictor $\tilde{g}$:
@@
16:  Use EK-FAC method to accelerate iHVP problem for $\hat{f}$
17:  Denote the pre-activated output of $\hat{f}$ as $f^{\prime}$, Compute $A$ as
$A=\frac{1}{n}\cdot\sum_{i=1}^{n}\hat{g}(x_{i})\cdot\hat{g}(x_{i})^{\mathrm{T}}$
18:  Comput $B$ as:
19:  Perform eigenvalue decomposition of $A$ and $B$, obtain $Q_{A},{\Lambda}_{{A}},{Q}_{B},{\Lambda}_{B}$, which satisfies
$A$ $=Q_{A}{\Lambda}_{{A}}{Q}_{{A}}^{\top}$
$B$ $={Q}_{B}{\Lambda}_{B}{Q}_{B}^{\top}$
20:  Define a diagonal matrix $\Lambda$ and compute the diagonal element as
21:  Compute $\hat{H}_{\hat{f}}^{-1}$ as
22:  Update label predictor $\tilde{f}$:
@@
23:  Return: $\tilde{f}$, $\tilde{g}$.
@@
3:  Use EK-FAC method to accelerate iHVP problem for $\hat{g}$
4:  for the $l$-th convolution layer of $\hat{g}$do
5:     Define the input activations $\{a_{j,t}\}$, weights $W=\left(w_{i,j,\delta}\right)$, and biases $b=\left(b_{i}\right)$ of this layer;
6:     Obtain the expanded activations $\llbracket{{A}}_{l-1}\rrbracket$ as:
7:     Compute the pre-activations:
8:     During the backpropagation process, obtain the $\mathcal{D}s_{i,t}$ as:
$\mathcal{D}s_{i,t}=\frac{\partial\sum_{j\notin M}^{k}\sum_{i=1}^{n}L_{C_{j}}}{\partial s_{i,t}}$
9:     Compute $\hat{\Omega}_{l-1}$ and $\hat{\Gamma_{l}}$:
10:     Perform eigenvalue decomposition of $\hat{\Omega}_{l-1}$ and $\hat{\Gamma}_{l}$, obtain $Q_{\Omega},{\Lambda}_{{\Omega}},{Q}_{\Gamma},{\Lambda}_{\Gamma}$, which satisfies
$\hat{\Omega}_{l-1}$ $=Q_{\Omega}{\Lambda}_{{\Omega}}{Q}_{{\Omega}}^{\top}$
$\hat{\Gamma}_{l}$ $={Q}_{\Gamma}{\Lambda}_{\Gamma}{Q}_{\Gamma}^{\top}$
11:     Define a diagonal matrix $\Lambda$ and compute the diagonal element as
12:     Compute $\hat{H}_{l}^{-1}$ as
13:  end for
14:  Splice $H_{l}$ sequentially into large diagonal matrices
15:  Update concept predictor $\tilde{g}$:
@@
16:  Use EK-FAC method to accelerate iHVP problem for $\hat{f}$
17:  Denote the pre-activated output of $\hat{f}$ as $f^{\prime}$, Compute $A$ as
$A=\frac{1}{n}\cdot\sum_{i=1}^{n}\hat{g}(x_{i})\cdot\hat{g}(x_{i})^{\mathrm{T}}$
18:  Comput $B$ as:
19:  Perform eigenvalue decomposition of $A$ and $B$, obtain $Q_{A},{\Lambda}_{{A}},{Q}_{B},{\Lambda}_{B}$, which satisfies
$A$ $=Q_{A}{\Lambda}_{{A}}{Q}_{{A}}^{\top}$
$B$ $={Q}_{B}{\Lambda}_{B}{Q}_{B}^{\top}$
20:  Define a diagonal matrix $\Lambda$ and compute the diagonal element as
21:  Compute ${H}_{\hat{f}}^{-1}$ as
${H}_{\hat{f}}^{-1}=\left(Q_{A}\otimes Q_{B}\right)\left(\Lambda+\lambda I_{d}\right)^{-1}\left(Q_{A}\otimes Q_{B}\right)^{\mathrm{T}}$
22:  Update label predictor $\tilde{f}$:
@@
23:  Map $\tilde{g}^{*}$ to $\tilde{g}$ by removing the $r$-th row of the matrix in the final layer of $\tilde{g}^{*}$ for $r\in M$.
24:  Return: $\tilde{f}$, $\tilde{g}$.
@@
3:  Use EK-FAC method to accelerate iHVP problem for $\hat{g}$
4:  for the ll-th convolution layer of $\hat{g}$do
5:     Define the input activations $\{a_{j,t}\}$, weights $W=\left(w_{i,j,\delta}\right)$, and biases $b=\left(b_{i}\right)$ of this layer;
6:     Obtain the expanded activations $\llbracket{{A}}_{l-1}\rrbracket$ as:
7:     Compute the pre-activations:
8:     During the backpropagation process, obtain the $\mathcal{D}s_{i,t}$ as:
9:     Compute $\hat{\Omega}_{l-1}$ and $\hat{\Gamma_{l}}$:
10:     Perform eigenvalue decomposition of $\hat{\Omega}_{l-1}$ and $\hat{\Gamma}_{l}$, obtain $Q_{\Omega},{\Lambda}_{{\Omega}},{Q}_{\Gamma},{\Lambda}_{\Gamma}$, which satisfies
$\hat{\Omega}_{l-1}$ $=Q_{\Omega}{\Lambda}_{{\Omega}}{Q}_{{\Omega}}^{\top}$
$\hat{\Gamma}_{l}$ $={Q}_{\Gamma}{\Lambda}_{\Gamma}{Q}_{\Gamma}^{\top}$
11:     Define a diagonal matrix $\Lambda$ and compute the diagonal element as
12:     Compute $\hat{H}_{l}^{-1}$ as
13:  end for
14:  Splice $H_{l}$ sequentially into large diagonal matrices
${H}_{\hat{g}}^{-1}=\left(\begin{array}[]{ccc}\hat{H}_{1}^{-1}&&\mathbf{0}\\ &\ddots&\\ \mathbf{0}&&\hat{H}_{d}^{-1}\end{array}\right)$
15:  Update concept predictor $\tilde{g}$:
@@
16:  Use EK-FAC method to accelerate iHVP problem for $\hat{f}$
17:  Denote the pre-activated output of $\hat{f}$ as $f^{\prime}$, Compute $A$ as
$A=\frac{1}{n}\cdot\sum_{i=1}^{n}\hat{g}(x_{i})\cdot\hat{g}(x_{i})^{\mathrm{T}}$
18:  Comput $B$ as:
19:  Perform eigenvalue decomposition of $A$ and $B$, obtain $Q_{A},{\Lambda}_{{A}},{Q}_{B},{\Lambda}_{B}$, which satisfies
$A$ $=Q_{A}{\Lambda}_{{A}}{Q}_{{A}}^{\top}$
$B$ $={Q}_{B}{\Lambda}_{B}{Q}_{B}^{\top}$
20:  Define a diagonal matrix $\Lambda$ and compute the diagonal element as
21:  Compute ${H}_{\hat{f}}^{-1}$ as
${H}_{\hat{f}}^{-1}=\left(Q_{A}\otimes Q_{B}\right)\left(\Lambda+\lambda I_{d}\right)^{-1}\left(Q_{A}\otimes Q_{B}\right)^{\mathrm{T}}$
Compute $A$ as:
@@
22:  Use EK-FAC method to accelerate iHVP problem for $\bar{f}$
23:  Denote the pre-activated output of $\bar{f}$ as $f^{\prime}$, Compute $A$ as
$A=\frac{1}{n}\cdot\sum_{i\in[n]-G}\hat{g}(x_{i})\cdot\hat{g}(x_{i})^{\mathrm{T}}$
24:  Comput $B$ as:
$B=\frac{1}{n}\cdot\sum_{i\in[n]-G}\nabla_{f^{\prime}}L_{Y}(\bar{f}\left(\hat{g}(x_{i})\right),y_{i})\cdot{\nabla_{f^{\prime}}L_{Y}(\bar{f}\left(\hat{g}(x_{i})\right),y_{i})}^{\mathrm{T}}$
25:  Perform eigenvalue decomposition of $A$ and $B$, obtain $Q_{A},{\Lambda}_{{A}},{Q}_{B},{\Lambda}_{B}$, which satisfies
$A$ $=Q_{A}{\Lambda}_{{A}}{Q}_{{A}}^{\top}$
$B$ $={Q}_{B}{\Lambda}_{B}{Q}_{B}^{\top}$
26:  Define a diagonal matrix $\Lambda$ and compute the diagonal element as
$\Lambda_{ii}^{*}=n^{-1}\sum_{j=1}^{n}\left(\left(Q_{A}\otimes Q_{B}\right)\nabla_{\bar{f}}L_{Y_{j}}\right)^{2}_{i}.$
27:  Compute ${H}_{\bar{f}}^{-1}$ as
${H}_{\bar{f}}^{-1}=\left(Q_{A}\otimes Q_{B}\right)\left(\Lambda+\lambda I_{d}\right)^{-1}\left(Q_{A}\otimes Q_{B}\right)^{\mathrm{T}}$
Compute $B^{\prime}$ as
@@
28:  Return: $\tilde{f}$, $\tilde{g}$.
Appendix F Additional Experiments
F.1 Experimental Setting
Dataset. We utilize two datasets: X-ray grading (OAI) [29] and Bird identification (CUB) [35]. X-ray grading (OAI) is a multi-center observational study of knee osteoarthritis, comprises 36,369 data points, and evaluates our model’s capability in medical imaging contexts. Specifically, we configure n=10 concepts that characterize crucial osteoarthritis indicators such as joint space narrowing, osteophytes, and calcification. Bird identification (CUB) consists of 11,788 data points, allowing us to assess our model in a fine-grained visual categorization task. Following the same setting outlined in [25], we employ 112 binary attributes to describe detailed visual features of birds, such as wing color, beak shape, and under-tail colors.
F.2 More Visualization Results
We provide our additional visualization results in Figure 5 and 6.
@@ F.2 More Visualization Results
Figure 5: Visualization of the top-10 most influential concepts for different classes.
@@
Figure 6: Visualization of the top-10 most influential concepts for different classes.
Appendix G More Related Work

### v2

@@ B.1 EK-FAC for CBMs
Besides, George et al. (2018) proposed a new method that corrects the error in equation 14 which sets the $i$-th diagonal element of $\Lambda_{\Omega_{l-1}}\otimes\Lambda_{\Gamma_{l}}$ as
@@
Appendix C Proof of Concept-label-level Influence
@@
Because of equation 15, the first term of equation C equals 00. Then we have
@@
where $H_{\hat{f}}=\nabla^{2}_{\hat{f}}\sum_{i=1}^{n}L_{Y_{i}}(\hat{f},\hat{g})$ is the Hessian matrix of the loss function respect to $\hat{f}$, $L_{Y_{i}}(\hat{f},\hat{g})\triangleq L_{Y}(\hat{f}(\hat{g}(x_{i})),y_{i})$, and $\bar{g}_{e}$ is given in Theorem C.1.
@@
Appendix D Proof of Concept-level Influence
@@
The main ideas are as follows: (i) First, we define a new predictor $\hat{g}^{*}_{p_{M}}$, which has the same dimension as $\hat{g}$ and the same output as $\hat{g}_{-p_{M}}$. Then deduce an approximation for $\hat{g}^{*}_{p_{M}}$. (ii) Then, we consider setting $p_{r}=0$ instead of removing it, we get $\hat{f}_{p_{M}=0}$, which is equivalent to $\hat{f}_{-p_{M}}$ according to lemma D.1. We estimate this new predictor as a substitute. (iii) Next, we assume we only use the updated concept predictor $\hat{g}_{p_{M}}^{*}$ for one data $(x_{i_{r}},y_{i_{r}},c_{i_{r}})$ and obtain a new label predictor $\hat{f}_{ir}$, and obtain a one-step Newtonian iterative approximation of $\hat{f}_{ir}$ with respect to $\hat{f}$. (iv) Finally, we repeat the above process for all data points and combine the estimate of $\hat{g}$ in Theorem D.2, we obtain a closed-form solution of the influence function for $\hat{f}$.
@@
Now, we come to the approximation of $\hat{f}_{-p_{M}}$. Noticing that the input dimension of $f$ decreases to $k-|M|$. We consider setting $p_{r}=0$ for all data points in the training phase of the label predictor and get another optimal model $\hat{f}_{p_{M}=0}$. From lemma D.1, we know that for the same input $x$, $\hat{f}_{p_{M}=0}(x)=\hat{f}_{-p_{M}}$. And the values of the corresponding parameters in $\hat{f}_{p_{M}=0}$ and $\hat{f}_{-p_{M}}$ are equal.
@@
Theorem D.2 gives us the edited version of $\hat{g}^{*}_{-p_{M}}$. Substitute it into equation 25, and we get the final closed-form edited label predictor under concept level:
@@
Appendix E Proof of Data-level Influence
@@
The edited version of $\hat{g}_{-z_{G}}$ has been deduced as $\bar{g}_{-z_{G}}$ in theorem E.1, substituting this approximation into equation 37, then we have
@@
Combining equation 34 and equation 39, then we deduce the final closed-form edited label predictor as
@@ Appendix F Algorithm
  Algorithm 4 EK-FAC for Concept Predictor $g$
1:  Input: Dataset $\mathcal{D}=\{(x_{i},y_{i},c_{i})\}_{i=1}^{N}$, original concept predictor $\hat{g}$.
2:  for the $l$-th convolution layer of $\hat{g}$do
3:     Define the input activations $\{a_{j,t}\}$, weights $W=\left(w_{i,j,\delta}\right)$, and biases $b=\left(b_{i}\right)$ of this layer;
4:     Obtain the expanded activations $\llbracket{{A}}_{l-1}\rrbracket$ as:
5:     Compute the pre-activations:
6:     During the backpropagation process, obtain the $\mathcal{D}s_{i,t}$ as:
7:     Compute $\hat{\Omega}_{l-1}$ and $\hat{\Gamma_{l}}$:
8:     Perform eigenvalue decomposition of $\hat{\Omega}_{l-1}$ and $\hat{\Gamma}_{l}$, obtain $Q_{\Omega},{\Lambda}_{{\Omega}},{Q}_{\Gamma},{\Lambda}_{\Gamma}$, which satisfies
$\hat{\Omega}_{l-1}$ $=Q_{\Omega}{\Lambda}_{{\Omega}}{Q}_{{\Omega}}^{\top}$
$\hat{\Gamma}_{l}$ $={Q}_{\Gamma}{\Lambda}_{\Gamma}{Q}_{\Gamma}^{\top}$
9:     Define a diagonal matrix $\Lambda$ and compute the diagonal element as
10:     Compute $\hat{H}_{l}^{-1}$ as
11:  end for
12:  Splice $H_{l}$ sequentially into large diagonal matrices
13:  Return: the inverse Hessian matrix $\hat{H}_{\hat{g}}^{-1}$.
  Algorithm 5 EK-FAC for Label Predictor $f$
1:  Input: Dataset $\mathcal{D}=\{(x_{i},y_{i},c_{i})\}_{i=1}^{N}$, original label predictor $\hat{f}$.
2:  Denote the pre-activated output of $\hat{f}$ as $f^{\prime}$, Compute $A$ as
$A=\frac{1}{n}\cdot\sum_{i=1}^{n}\hat{g}(x_{i})\cdot\hat{g}(x_{i})^{\mathrm{T}}$
3:  Comput $B$ as:
4:  Perform eigenvalue decomposition of AA and BB, obtain $Q_{A},{\Lambda}_{{A}},{Q}_{B},{\Lambda}_{B}$, which satisfies
$A$ $=Q_{A}{\Lambda}_{{A}}{Q}_{{A}}^{\top}$
$B$ $={Q}_{B}{\Lambda}_{B}{Q}_{B}^{\top}$
5:  Define a diagonal matrix $\Lambda$ and compute the diagonal element as
6:  Compute $\hat{H}_{\hat{f}}^{-1}$ as
7:  Return: the inverse Hessian matrix $\hat{H}_{\hat{f}}^{-1}$.
@@
3:  Use EK-FAC method in algorithm F to accelerate iHVP problem for $\hat{g}$ and obtain the inverse Hessian matrix $\hat{H}_{\hat{g}}^{-1}$
4:  Update concept predictor $\tilde{g}$:
@@
5:  Use EK-FAC method in algorithm F to accelerate iHVP problem for $\hat{f}$ and obtain $\hat{H}_{\hat{f}}^{-1}$
6:  Update label predictor $\tilde{f}$:
@@
7:  Return: $\tilde{f}$, $\tilde{g}$.
@@
3:  Use EK-FAC method in algorithm F to accelerate iHVP problem for $\hat{g}$ and obtain the inverse Hessian matrix $\hat{H}_{\hat{g}}^{-1}$
4:  Update concept predictor $\tilde{g}$:
@@
5:  Use EK-FAC method in algorithm F to accelerate iHVP problem for $\hat{f}$ and obtain $\hat{H}_{\hat{f}}^{-1}$
6:  Update label predictor $\tilde{f}$:
@@
8:  Return: $\tilde{f}$, $\tilde{g}$.
@@
3:  Use EK-FAC method in algorithm F to accelerate iHVP problem for $\hat{g}$ and obtain the inverse Hessian matrix $\hat{H}_{\hat{g}}^{-1}$
4:  Update concept predictor $\tilde{g}$:
@@
5:  Use EK-FAC method in algorithm F to accelerate iHVP problem for $\hat{f}$ and obtain $H_{\hat{f}}^{-1}$ Compute $A$ as:
@@
6:  Use EK-FAC method in algorithm F to accelerate iHVP problem for $\bar{f}$ and obtain ${H}_{\bar{f}}^{-1}$ Compute $B^{\prime}$ as
@@
7:  Return: $\tilde{f}$, $\tilde{g}$.
Appendix G Additional Experiments
G.1 Experimental Setting
RMIA score. The RMIA score is computed as:
$LR_{\theta}(x,z)\approx\frac{\text{Pr}(f_{\theta}(x)|\mathcal{N}(\mu_{x,\bar{z}}(x),\sigma^{2}_{x,\bar{z}}(x)))}{\text{Pr}(f_{\theta}(x)|\mathcal{N}(\mu_{\bar{x},z}(x),\sigma^{2}_{\bar{x},z}(x)))}\times\frac{\text{Pr}(f_{\theta}(z)|\mathcal{N}(\mu_{x,\bar{z}}(z),\sigma^{2}_{x,\bar{z}}(z)))}{\text{Pr}(f_{\theta}(z)|\mathcal{N}(\mu_{\bar{x},z}(z),\sigma^{2}_{\bar{x},z}(z)))}$
where $f_{\theta}(x)$ represents the model’s output (logits) for the data point $x$, $\mathcal{N}(\mu,\sigma^{2})$ denotes a Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$, $\mu_{x,\bar{z}}(x)$ is the mean of the model’s outputs for $x$ under the assumption that $x$ belongs to the training set, and $\sigma^{2}_{x,\bar{z}}(x)$ is the variance of the model’s outputs for $x$. The likelihoods $\text{Pr}(f_{\theta}(x)|\mathcal{N})$ represent the probability that the model’s output $f_{\theta}(x)$ follows the Gaussian distribution parameterized by $\mu$ and $\sigma^{2}$, under the two different hypotheses: $x$ being a member of the training set versus not being a member.
G.2 More Visualization Results
We provide our additional visualization results in Figure 6, 7, 8, 9, 10, 11, and 12.
Figure 6: Visualization of the top-10 most influential concepts for different classes in CUB.
Figure 7: Visualization of the top-10 most influential concepts for different classes in CUB.
@@ G.2 More Visualization Results
Figure 8: Visualization of the most influential concept label related to different data in CUB.
@@
Figure 9: Visualization of the most influential concept label related to different data in CUB.
Figure 10: Visualization of the most influential concept label related to different data in CUB.
Figure 11: Visualization of the most influential concept label related to different data in CUB.
Figure 12: Visualization of the most influential concept label related to different data in OAI.
Appendix H More Related Work